### v2

# arXiv:2601.21236v2[cs.IT] 02 Mar 2026

Abstract
We study deletion-correcting codes for an adversarial nanopore channel in which at most $t$ deletions may occur. We propose an explicit construction of $q$-ary codes of length $n$ for this channel with $2t\log_{q}n+\Theta(\log\log n)$ redundant symbols. We also show that the optimal redundancy is between $t\log_{q}n+\Omega(1)$ and $2t\log_{q}n-\log_{q}\log_{2}n+O(1)$, so our explicit construction matches the existential upper bound to first order. In contrast, for the classical adversarial $q$-ary deletion channel, the smallest redundancy achieved by known explicit constructions that correct up to $t$ deletions is $4t(1+\epsilon)\log_{q}n+o(\log n)$.
I Introduction
Nanopore sequencing has attracted considerable attention for DNA-based data storage due to its advantages over other sequencing technologies. However, it is prone to several types of errors, including deletion, duplication and substitution errors. In essence, nanopore sequencing can be modeled as passing a string through an inter-symbol interference (ISI) channel, followed by additional noisy channels that further corrupt the ISI channel output. Specifically, the ISI channel scans the string from left to right using a sliding window of size $\ell$, shifted by one position each step, thereby producing a sequence of $\ell$-mers. The subsequent noisy channels then introduce errors into this $\ell$-mer sequence. The goal of nanopore sequencing is to recover the original string from the final noisy output.
A large body of work has studied nanopore sequencing in recent years. Depending on whether the underlying channel model is probabilistic, the literature can be grouped into two broad classes. Under a probabilistic channel model, Magner et al. [1] gave an information-theoretic analysis of the impact of sticky insertion-deletion errors in nanopore sequencers, without explicitly accounting for ISI. Later, Mao et al. [2] developed a signal-level model for nanopore sequencing that captures ISI and analyzed achievable information rates. In [3, 4, 5], the authors studied a nanopore channel model based on a Markov chain with memory of order $\ell$. In subsequent work, McBain et al. [6, 7] introduced a channel model in the language of semi-Markov processes and estimated its achievable information rates using efficient algorithms. The achievable information rates of this model were further examined in [8, 9]. A number of other works have also considered the design of coding schemes for nanopore sequencing under various probabilistic models. We refer the interested reader to the recent works [10, 11, 12, 13, 14] and the references therein.
Another line of research assumes deterministic channel models and investigates the combinatorial and algebraic coding problems arising in nanopore sequencing. Specifically, the authors of [15] introduced a deterministic model capturing only ISI and substitution errors, and proposed coding schemes for error correction. Extending the work of [15], the authors of [16] studied coding schemes that can deal with duplication and deletion errors. In addition, a number of recent papers model the noisy output of deterministic nanopore channels as a sequence of erroneous $\ell$-mer compositions. Under this framework, the work [17] considered ISI effect with substitution errors and gave a coding scheme that can correct a single substitution with near-optimal redundancy. Extensions to multiple substitution errors were further explored in [18, 19]. Assuming instead that only deletion errors are present in addition to ISI, the work [20] proposed a single-deletion-correcting code with near-optimal redundancy. Meanwhile, the paper [21] studied ISI with tandem-duplication errors in the $\ell$-mer composition sequence and analyzed the optimal code rate. We also note that the recent works [22, 23] also studied various coding-theoretic questions under the assumption that the noisy channel output is in the form of $\ell$-mer compositions. Lastly, motivated by the fact that basecallers are often used to infer a sequence of $\ell$-mers from the output of practical nanopore sequencers, the work of [24] modeled the noisy channel output as a sequence of erroneous $\ell$-mers and constructed codes that can correct up to two $\ell$-mer deletions.
In this work, under the same assumption as in [24], we study bounds and constructions for codes that can correct a constant number of deletions. We note that modeling the noisy channel output as a sequence of $\ell$-mers yields a model that resembles the $\ell$-symbol read channel [25, 26, 27]. However, the reads in the $\ell$-symbol read channel are taken cyclically. That is, after reading the last $\ell$ symbols of the input, the channel continues with the size-$\ell$ window that wraps around the end of the string, ending with the window consisting of the last symbols followed by the first $\ell-1$ symbols. In [27], Chee et al. introduced locally-constrained de Bruijn codes for correcting synchronization errors in $\ell$-symbol read channels. For the case of $\ell=2$, codes correcting deletion and substitution errors were studied in [28].
We note that the nanopore channel considered in this work is also closely related to the classical adversarial deletion channel [29]. For binary codes, Levenshtein [29] showed that the optimal redundancy for $t$-deletion-correcting binary codes of length $n$ lies between $t\log_{2}n+o(\log n)$ and $2t\log_{2}n+o(\log n)$, where $t$ is a constant independent of $n$. For a single deletion, the Varshamov-Tenengolts construction [29, 30] is known to have redundancy at most $\log_{2}(n+1)$, and for two deletions, Guruswami and Håstad [31] constructed codes with redundancy $4\log_{2}n+O(\log\log n)$, matching the existential upper bound. More generally, explicit constructions of $t$-deletion-correcting binary codes with order-optimal redundancy are presented in [32, 33]. In particular, the construction of Sima and Bruck [32] achieves redundancy $8t\log_{2}n+o(\log n)$. For $q$-ary codes, Levenshtein [34] proved analogous redundancy bounds $t\log_{q}n+o(\log n)$ and $2t\log_{q}n+o(\log n)$ (measured in $q$-ary symbols). A construction with nearly optimal redundancy for a single deletion was given by Tenengolts [35]. In [36], Sima et al. proposed $t$-deletion-correcting $q$-ary codes that require at most $4t(1+\epsilon)\log_{q}n+o(\log n)$ redundant symbols. On the existential side, Alon et al. [37] obtained the first asymptotic improvement to Levenshtein’s lower bound on the optimal code size, while Levenshtein’s upper bound has been improved in various parameter regimes in a series of works [38, 39, 40, 41, 42, 43].
Our contributions
We give an explicit construction of $q$-ary codes of length $n$ with $2t\log_{q}n+\Theta(\log\log n)$ redundant symbols that can correct up to $t$ deletions of $\ell$-mers, where $t\leq\min\{\ell-2,(\ell+1)/2\}$ is a constant independent of $n$. Our construction reduces deletion correction to an almost Hamming-type error-correction problem, and uses generalized Reed-Solomon (GRS) codes as a key building block. In addition, we observe that our nanopore channel model can be viewed as a weaker version of the classical adversarial deletion channel. Consequently, the lower bound of Alon et al. applies directly and yields an existential upper bound $2t\log_{q}n-\log_{q}\log_{2}n+O(1)$ on the optimal redundancy. This further suggests our explicit construction matches this existential upper bound to first order. Moreover, drawing on ideas of Levenshtein [29], we use the probabilistic method to establish an upper bound on the optimal code size, implying a lower bound of $t\log_{q}n+\Omega(1)$ on the optimal redundancy.
Paper outline
The rest of the paper is organized as follows. Section II formally introduces our channel model and present a few technical lemmas used throughout the paper. Section III is devoted to the explicit construction of $t$-deletion-correcting codes for our channel model. Section IV presents existential bounds on the optimal code size. Finally, Section V discusses open problems and future directions.
II Channel model and technical lemmas
[…]
The central question studied in this work is the recovery of the $q$-ary string $\mathbf{x}$ from the sequence of $\ell$-mers $\mathbf{z}_{0}^{{n+\ell}-\tau}$. Since any two consecutive $\ell$-mers in $\mathbf{s}_{0}^{n+\ell}$ share $\ell-1$ symbols due to the ISI effect, a natural first step for reconstructing $\mathbf{x}$ is to recover some (if not all) of the deleted $\ell$-mers from the output $\ell$-mers $\mathbf{z}_{0}^{n+\ell-\tau}$ by utilizing the shared symbols. To this end, let us introduce the concept of consistent $\ell$-mers. We call two consecutive $\ell$-mers $\mathbf{z}_{i}$ and $\mathbf{z}_{i+1}$ consistent if the last $\ell-1$ symbols of $\mathbf{z}_{i}$ are the same as the first $\ell-1$ symbols of $\mathbf{z}_{i+1}$. Otherwise, we say that $\mathbf{z}_{i}$ and $\mathbf{z}_{i+1}$ are inconsistent and there is inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$. It is clear that any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$ must result from deletions introduced by the channel. Furthermore, since any pattern of deletions can be viewed as a collection of disjoint bursts of deletions, [^2] In this paper, a deletion burst always refers to a maximal deletion burst. any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$ must be attributed to some deletion burst. By [24, Lemma 1], which we restate below, if there is any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$ caused by bursts of at most $\ell-2$ deletions, then we can recover some (if not all) of the deleted $\ell$-mers so that the inconsistency is eliminated, by looking at the output sequence of $\ell$-mers.
Lemma 1 ([24, Lemma 1]).
Assume $\ell>2$ and there are deletion bursts of length at most $\ell-2$ in $\mathbf{z}_{0}^{{n+\ell}-\tau}$. If $\mathbf{z}_{i}$ and $\mathbf{z}_{i+1}$ are inconsistent, then at least one $\ell$-mer has been deleted between $\mathbf{s}_{\zeta(i)}$ and $\mathbf{s}_{\zeta(i+1)}$. Moreover, a minimum number of deleted $\ell$-mers can be recovered from $\mathbf{z}_{i},\mathbf{z}_{i+1}$, and inserted between $\mathbf{z}_{i},\mathbf{z}_{i+1}$ to form a sequence in which any two consecutive $\ell$-mers are consistent.
Therefore, given the channel output $\mathbf{z}_{0}^{{n+\ell}-\tau}$, we can remove any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$, thereby obtaining a sequence of consistent $\ell$-mers. In other words, deletions that give rise to inconsistent $\ell$-mers are generally easy to correct. In view of this, we henceforth assume that the channel output $\mathbf{z}_{0}^{{n+\ell}-\tau}$ is already a sequence of consistent $\ell$-mers. With ensured consistency, we can then transform $\mathbf{z}_{0}^{{n+\ell}-\tau}$ into a $q$-ary string $\tilde{\mathbf{y}}$ of length $n+2l-\tau$ by concatenating all the symbols of the first $\ell$-mer $\mathbf{z}_{0}$ with the last symbol of each subsequent $\ell$-mers $\mathbf{z}_{1},\ldots,\mathbf{z}_{{n+\ell}-\tau}$. For simplicity, we denote the mapping from the output $\ell$-mers $\mathbf{z}_{0}^{{n+\ell}-\tau}\in(\Sigma_{q}^{\ell})^{{n+\ell}-\tau}$ to $\tilde{\mathbf{y}}\in\Sigma_{q}^{{n+2\ell}-\tau}$ by $\psi$ and refer to $\tilde{\mathbf{y}}=\psi(\mathbf{z}_{0}^{{n+\ell}-\tau})$ as the output string of the nanopore channel corresponding to the input $\tilde{\mathbf{x}}$. Since the mapping $\psi$ is injective, we have the following simple proposition.
Proposition 2.
A code $\mathscr{C}\subset\Sigma_{q}^{n}$ can recover $\mathbf{x}\in\mathscr{C}$ from the output $\ell$-mers $\mathbf{z}_{0}^{n+\ell-\tau}$ corresponding to the input string $\tilde{\mathbf{x}}$ if and only if it can recover $\mathbf{x}$ from the output string $\tilde{\mathbf{y}}$.
Note that Lemma 1 only guarantees the recovery of a minimum number of deleted $\ell$-mers necessary to restore consistency. However, there may be $\ell$-mers deleted in $\mathbf{z}_{0}^{{n+\ell}-\tau}$ whose removal does not cause any inconsistency. To address this, we next characterize substrings of $\tilde{\mathbf{x}}$ that lead to such $\ell$-mers, i.e., $\ell$-mers that preserve consistency even when some of them are deleted. It turns out that these substrings inherently possess a periodic structure.
Definition 1 (Periodic patterns).
A string $\mathbf{s}=(s_{1},\ldots,s_{N})\in\Sigma_{q}^{N}$ is said to have period $T\leq N-1$ if $s_{i}=s_{i+T}$ for all $i=1,\ldots,N-T$. The least period of $\mathbf{s}$ is the smallest positive integer $T$ with this property. A substring of $\mathbf{s}$} is called a periodic pattern with period $T$ if the substring has period $T$. A periodic pattern with period $T$ in $\mathbf{s}$ is called maximal if it cannot be extended to the left or right within $\mathbf{s}$ to form a longer substring that has the same period $T$.
The following theorem of Fine and Wilf [44] is helpful for our subsequent discussion on periodic patterns.
Theorem 3 (Fine-Wilf theorem [44, Theorem 1]).
Let $(f_{i})_{i\geq 0},(g_{i})_{i\geq 0}$ be two strings over $\Sigma_{q}$ with period $T_{1},T_{2}$, respectively. If $f_{i}=g_{i}$ for $T_{1}+T_{2}-\gcd(T_{1},T_{2})$ consecutive integer $i$, then $f_{i}=g_{i}$ for all $i$. The result would be false if $T_{1}+T_{2}-\gcd(T_{1},T_{2})$ were replaced by anything smaller.
Remark II.1.
One direct implication of Theorem 3 is that, within a string, two adjacent maximal periodic patterns with period $T_{1},T_{2}$ can overlap in at most $T_{1}+T_{2}-\gcd(T_{1},T_{2})-1$ positions. Note that if the length of one maximal periodic pattern is at most $T_{1}+T_{2}-\gcd(T_{1},T_{2})-1$, it is possible that one pattern is contained in another.
Remark II.2.
If a periodic pattern with period $T$ has length at least $2T$, then its least period must divide $T$. This is because if $T^{\prime}\leq T$ is the least period, then $2T\geq T+T^{\prime}-\gcd(T,T^{\prime})$. Therefore, by Theorem 3, the pattern also has period $\gcd(T,T^{\prime})$, which divides $T^{\prime}$. But $T^{\prime}$ is the least period so $\gcd(T,T^{\prime})=T^{\prime}$ and $T^{\prime}|T$.
Generalizing [24, Lemma 2], the following lemma gives a necessary and sufficient condition, in terms of periodic patterns, under which a single burst of deletions of $\ell$-mers does not introduce any inconsistency. Since any number of deletions can be viewed as a collection of multiple bursts of deletions, this yields a characterization of deletions that preserve consistency in terms of periodic patterns.
Lemma 4.
Assume $t\leq\ell-2$ and that there is a single deletion burst of length $\tau\leq t$ between $\mathbf{s}_{i}$ and $\mathbf{s}_{i+\tau+1}$, i.e., $\mathbf{s}_{i+1},\dots,\mathbf{s}_{i+\tau}$ are deleted but $\mathbf{s}_{i}$ and $\mathbf{s}_{i+\tau+1}$ are not. Then the burst of deletions does not result in any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$ if and only if $(x_{i-\ell+2},\dots,x_{i+\tau})$ is a length-$(\tau+\ell-1)$ periodic pattern whose least period divides $\tau$.
Proof.
Since $\mathbf{s}_{i+1},\dots,\mathbf{s}_{i+\tau}$ are deleted but $\mathbf{s}_{i}$ and $\mathbf{s}_{i+\tau+1}$ are not, there exist $i^{*}$ such that $\mathbf{z}_{i^{*}}=\mathbf{s}_{i}$ and $\mathbf{z}_{i^{*}+1}=\mathbf{s}_{i+\tau+1}$, i.e., $i=\zeta(i^{*})$ and $i+\tau+1=\zeta(i^{*}+1)$.
First, assume the deletion burst does not cause any inconsistency in $\mathbf{z}_{0}^{{n+\ell}-\tau}$. Then the last $\ell-1$ symbols of $\mathbf{z}_{i^{\star}}=(x_{\zeta(i^{\star})-\ell+1},\ldots,x_{\zeta(i^{*})})$ and the first $\ell-1$ symbols of $\mathbf{z}_{i^{\star}+1}=(x_{\zeta(i^{*}+1)-\ell+1},\ldots,x_{\zeta(i^{*}+1)})$ are the same. Therefore, we have $(x_{i-\ell+2},\dots,x_{i})=(x_{\zeta(i^{\star})-\ell+2},\dots,x_{\zeta(i^{\star})})=(x_{\zeta(i^{\star}+1)-\ell+1},\dots,x_{\zeta(i^{\star}+1)-1})=(x_{i+\tau-\ell+2},\dots,x_{i+\tau})$. Hence, we have $x_{j}=x_{j+\tau}$ for all $j=i-\ell+2,\dots,i$. Therefore, $(x_{i-\ell+2},\dots,x_{i+\tau})$ is a periodic pattern with period $\tau$ and length $\tau+\ell-1$. Since $t\leq\ell-2$, we have $\tau+\ell-1>2\tau$, and thus by Remark II.2 the least period of $(x_{i-\ell+2},\dots,x_{i+\tau})$ divides $\tau$.
Next, assume $(x_{i-\ell+2},\dots,x_{i+\tau})$ is a periodic pattern whose least period divides $\tau$. Therefore, $x_{j}=x_{j+\tau}$ for $j=i-\ell+2,\dots,i$. It follows that $(x_{\zeta(i^{\star})-\ell+2},\dots,x_{\zeta(i^{\star})})=(x_{i-\ell+2},\dots,x_{i})=(x_{i+\tau-\ell+2},\dots,x_{i+\tau})=(x_{\zeta(i^{\star}+1)-\ell+1},\dots,x_{\zeta(i^{\star}+1)-1})$. Hence, $\mathbf{z}_{i^{\star}}$ and $\mathbf{z}_{i^{\star}+1}$ are consistent. ∎
As a consequence of Lemma 4, the channel output $\mathbf{z}_{0}^{{n+\ell}-\tau}$ remains consistent if each deletion burst of $\ell$-mers arises in a periodic pattern in $\tilde{\mathbf{x}}$ whose length is at least $T+\ell-1$ where $T\leq t$ is the least period of the pattern. In view of this, we call such patterns consistency-preserving, formally defined below.
Definition 2 (Consistency-preserving pattern).
A substring of $\tilde{\mathbf{x}}$ is a consistency-preserving pattern if it is a maximal periodic pattern with least period $T\leq t$ and length at least $T+\ell-1$.
Remark II.3.
We make a few observations about consistency-preserving patterns. First, under the assumption $T\leq t\leq\ell-2$, a consistency-preserving pattern has length at least $2T$; that is, it contains at least two complete least periods. At the same time, since $T\geq 1$, the length of such a pattern is also at least $\ell$. In addition, Lemma 4 implies that a deletion burst of length $b$ in the output $\ell$-mers induces a deletion burst of the same length $b$ within a consistency-preserving pattern of the output string. If this consistency-preserving pattern has least period $T$, then it is necessary that $T|b$, so a deletion burst of length $b$ removes exactly $b/T$ complete periods of the pattern. Conversely, a consistency-preserving pattern of length $b+\ell-1$ with least period $T$ can admit deletion of at most $\lfloor b/T\rfloor$ complete least periods while keeping the output $\ell$-mers consistent.
III Explicit construction of codes
In this section we assume further that $t\leq(\ell+1)/2$ and present an explicit construction of codes $\mathscr{C}\subset\Sigma_{q}^{n}$ that can correct up to $t\leq\min\{\ell-2,(\ell+1)/2\}$ deletions of $\ell$-mers for the nanopore channel. Specifically, let $\mathbf{s}_{0},\mathbf{s}_{n+\ell}\in\Sigma_{q}^{\ell}$ be two fixed adapter sequences. For any input string $\tilde{\mathbf{x}}=(\mathbf{s}_{0},\mathbf{x},\mathbf{s}_{n+\ell})$ with $\mathbf{x}\in\mathscr{C}$, we show that one can recover $\mathbf{x}$ from the output $\ell$-mers $\mathbf{z}_{0}^{{n+\ell}-\tau}$ of the nanopore channel, provided that at most $t\leq\min\{\ell-2,(\ell+1)/2\}$ deletions occur in the output. Before delving into the technical details, below we outline the high-level idea of our construction.
As discussed earlier, the output $\ell$-mers are assumed to be consistent, so deletions only occur within consistency-preserving patterns. Note that if at least one complete least period in a consistency-preserving pattern survives the deletions in the output string, and the length of the pattern in the input string is known a priori, then the full pattern can be recovered by extending it using its periodicity. As mentioned in Remark II.3, for a consistency-preserving pattern of length $b+\ell-1$ with least period $T$, it can undergo deletion of at most $\lfloor b/T\rfloor$ complete least periods, leaving a substring of length $\ell+b-1-\lfloor b/T\rfloor T\geq\ell-1$. Thus, at least $\ell-1$ symbols of the consistency-preserving pattern survive the deletions. This also implies that at least one complete least period of the pattern necessarily remains, since $T\leq t\leq\ell-2$. Therefore, we need to keep track of the maximal periodic patterns of length at least $\ell-1$ in both the input string $\tilde{\mathbf{x}}$ and the output string $\tilde{\mathbf{y}}$, which we call check patterns.
Definition 3 (Check patterns).
A substring of $\mathbf{s}=(s_{1},\ldots,s_{N})\in\Sigma_{q}^{N}$ is a check pattern if it is a maximal periodic pattern with period at most $t$ and length at least $\ell-1$.
Now if the set of check patterns in the input string $\tilde{\mathbf{x}}$ is the same as in the output string $\tilde{\mathbf{y}}$, up to a reduction in the length of each pattern, then one can establish a one-to-one correspondence between check patterns in $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{y}}$ that matches each check pattern in $\tilde{\mathbf{x}}$ with the same pattern (possibly with reduced length) in $\tilde{\mathbf{y}}$. Thus, if the lengths of the check patterns in $\tilde{\mathbf{x}}$ are known, one can restore the check patterns in $\tilde{\mathbf{y}}$ to their original lengths using their periodicity. We now show that the check patterns present in the input string are preserved in the output string, possibly with reduced length. This follows from the technical lemma below.
Lemma 5.
Assume $t\leq\min\{\ell-2,(\ell+1)/2\}$. Any two adjacent check patterns in a string can overlap in at most $\ell-2$ positions.
Proof.
Since $t\leq(\ell+1)/2$, by Theorem 3 two adjacent check patterns with different period $T_{1},T_{2}$ can overlap in at most $T_{1}+T_{2}-\gcd(T_{1},T_{2})-1\leq 2t-3\leq\ell-2$ positions. Moreover, if $T_{1}=T_{2}$, then by Theorem 3 they can overlap in at most $T_{1}-1\leq t-1<\ell-2$ positions. Hence, any two adjacent check patterns in a string can overlap in at most $\ell-2$ positions. ∎
Lemma 6.
Assume $t\leq\min\{\ell-2,(\ell+1)/2\}$. The number of check patterns in $\tilde{\mathbf{x}}$ is the same as in $\tilde{\mathbf{y}}$. In particular, every check pattern in $\tilde{\mathbf{y}}$ arises from exactly one check pattern in $\tilde{\mathbf{x}}$, possibly with reduced length. Moreover, the check patterns in $\tilde{\mathbf{x}}$ start at distinct positions, and the same holds for $\tilde{\mathbf{y}}$.
Proof.
As mentioned above, deletion can only remove complete least periods of a consistency-preserving pattern in $\tilde{\mathbf{x}}$, leaving a periodic pattern of length at least $\ell-1$. Therefore, each check pattern in $\tilde{\mathbf{y}}$ must come from a check pattern in $\tilde{\mathbf{x}}$, and the number of check pattern in $\tilde{\mathbf{y}}$ is no larger than that in $\tilde{\mathbf{x}}$.
It remains to show that no two check patterns in $\tilde{\mathbf{x}}$ can be combined into a single check pattern in $\tilde{\mathbf{y}}$ through deletion. Note that by Lemma 5, any two adjacent check patterns in a string can overlap in at most $\ell-2$ positions. Since a check pattern has length at least $\ell-1$, this implies they cannot start (or end) at the same position, i.e., one pattern cannot be contained in another. For the same reason, this also implies any two adjacent check patterns in $\tilde{\mathbf{x}}$ remain to be distinct in $\tilde{\mathbf{y}}$. So we are left to show that two non-adjacent check patterns cannot be combined by deletion. For two non-adjacent check patterns in $\tilde{\mathbf{x}}$ to form a single check pattern in $\tilde{\mathbf{y}}$, the substring of $\tilde{\mathbf{x}}$ that separates them must be deleted. Moreover, this separating substring must lie within another check pattern in $\tilde{\mathbf{x}}$, which is adjacent to the two non-adjacent check patterns. By the argument above, this adjacent check pattern cannot be merged with either of the two, and thus any two non-adjacent check patterns remain distinct in $\tilde{\mathbf{y}}$. ∎
By Lemma 6, we can identify each check pattern in $\tilde{\mathbf{x}}$ with a unique check pattern in $\tilde{\mathbf{y}}$, sequentially from left to right. If the information of the lengths of the check patterns in $\tilde{\mathbf{x}}$ can be embedded in $\tilde{\mathbf{x}}$ and preserved in $\tilde{\mathbf{y}}$, then recovering $\tilde{\mathbf{x}}$ from $\tilde{\mathbf{y}}$ should be straightforward. The remaining task, therefore, is to design a method to encode these lengths in a way that is robust to deletions. To this end, we first examine how many check patterns in $\tilde{\mathbf{y}}$ may have reduced length.
Lemma 7.
Assume $t\leq\min\{\ell-2,(\ell+1)/2\}$. At most $\tau\leq t$ check patterns in $\tilde{\mathbf{y}}$ can have length smaller than their corresponding check patterns in $\tilde{\mathbf{x}}$.
Proof.
By Lemma 6, there is a one-to-one correspondence between check patterns in the input string $\tilde{\mathbf{x}}$ and the output string $\tilde{\mathbf{y}}$ that matches each check pattern in $\tilde{\mathbf{x}}$ with the same pattern (possibly with reduced length) in $\tilde{\mathbf{y}}$. Since there are $\tau$ deletions in the output $\ell$-mers, there are $\tau$ deleted symbols in the output string $\tilde{\mathbf{y}}$. Bursts are by definition disjoint, so there can be at most $\tau$ deletion bursts in $\tilde{\mathbf{y}}$, with equality if and only if each burst has length one. We next show that a single deletion burst can reduce the length of at most one check pattern in the input string $\tilde{\mathbf{x}}$.
Consider a deletion burst of length $b\leq\tau$ that deletes the symbols $(x_{i+1},\ldots,x_{i+b})$. By Lemma 4, the output $\ell$-mers remain consistent if and only if $(x_{i-\ell+2},\dots,x_{i+b})$ is a periodic pattern with least period $T$ satisfying $T\mid b$. Let $\mathbf{p}_{1}$ be the maximal periodic pattern in $\tilde{\mathbf{x}}$ with least period $T$ that contains $(x_{i-\ell+2},\dots,x_{i+b})$. It follows that $\mathbf{p}_{1}$ is a check pattern.
Suppose the same burst also shortens the length of another check pattern $\mathbf{p}_{2}$. Then the burst must overlap $\mathbf{p}_{2}$, and thus $\mathbf{p}_{1}$ overlaps $\mathbf{p}_{2}$. By Lemma 5, the size of the overlap between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ is at most $\ell-2$. Since check patterns have length at least $\ell-1$, it follows that neither $\mathbf{p}_{1}$ nor $\mathbf{p}_{2}$ is contained in the other, and thus the overlap between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ must be a suffix of one and a prefix of the other. Since the length of $\mathbf{p}_{1}$ is at least $b+\ell-1$, it follows that $\mathbf{p}_{1}$ contains at least $b+\ell-1-(\ell-2)>b$ consecutive symbols that are disjoint from $\mathbf{p}_{2}$.
Since $\mathbf{p}_{1}$ has period $T$ and $T\mid b$, deleting any $b$ consecutive symbols in $\mathbf{p}_{1}$ produces the same resulting string as deleting $(x_{i+1},\ldots,x_{i+b})$. Therefore, one may assume without loss of generality that the burst deletes any $b$ consecutive symbols of $\mathbf{p}_{1}$. In particular, one may replace the original burst by an equivalent burst of length $b$ that is disjoint from $\mathbf{p}_{2}$, without changing the output string. Under this equivalent burst, $\mathbf{p}_{2}$ is unaffected and cannot be shortened by this burst, which is a contradiction.
Therefore, a single deletion burst can shorten at most one check pattern. Since there are at most $\tau$ bursts, at most $\tau$ check patterns in $\tilde{\mathbf{y}}$ can have length smaller than their corresponding check patterns in $\tilde{\mathbf{x}}$. ∎
As a consequence of Lemma 7, if we record the lengths of the check patterns in $\tilde{\mathbf{x}}$ sequentially in one tuple and those in $\tilde{\mathbf{y}}$ in another, then the Hamming distance between these two tuples is at most $t$. Therefore, a Hamming-type error-correcting code with minimum distance at least $2t+1$ may be used to protect these lengths. To proceed, we shift our attention from $\tilde{\mathbf{x}}$ to $\bar{\mathbf{x}}:=(\mathbf{s}_{0},\mathbf{u})$, where $\mathbf{u}\in\Sigma_{q}^{k}$ is an information string. Note that the discussion above for $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{y}}$ applies equally to $\bar{\mathbf{x}}$ and its corresponding output string $\bar{\mathbf{y}}$ induced by the nanopore channel. By Lemma 6, the check patterns all start at distinct positions, and thus the number of check patterns in $\bar{\mathbf{x}}=(\mathbf{s}_{0},\mathbf{u})\in\Sigma_{q}^{k+\ell}$ is at most $k+1$. Our idea is to encode the tuple of check-pattern lengths of $\bar{\mathbf{x}}$ using a parity-check matrix of a GRS code of length $m:=k+1$ with distance $2t+1$ to obtain the syndrome of this tuple. This syndrome is then further encoded by a $q$-ary $t$-deletion-correcting code from [36]. Finally, the resulting encoding is appended to $\bar{\mathbf{x}}$, followed by $\mathbf{s}_{n+\ell}$, to form the full input string $\tilde{\mathbf{x}}$. We are now ready to describe the construction of our code $\mathscr{C}$.
III-A Encoding
Let $\phi\colon\Sigma_{q}^{\leq k+\ell}\to\{0,\ell-1,\ell,\ldots,k+\ell\}^{m}$ be a function that maps a $q$-ary string $\mathbf{s}$ of length at most $k+\ell$ to the $m$-tuple of its check pattern lengths, listed in order of their starting positions and padded with zeros if $\mathbf{s}$ contains fewer than $m$ check patterns. Note that each element in the $m$-tuple $\phi(\bar{\mathbf{x}})$ can take $k+3$ different values. As we would like to encode $\phi(\bar{\mathbf{x}})$ using a parity-check matrix of a length-$m$ GRS code, the field size of the code should be at least $\max\{k+3,m\}=k+3$.
Let $p$ be a prime power such that $p\geq k+3$. Let $f\colon\{0,\ell-1,\ell,\dots,k+\ell\}\to\mathbb{F}_{p}$ be an injective mapping. Furthermore, for a string $\mathbf{s}$ of length at most $k+\ell$, write $\phi(\mathbf{s})=(\phi(\mathbf{s})_{1},\ldots,\phi(\mathbf{s})_{m})$ and define $\Phi\colon\Sigma_{q}^{\leq k+\ell}\to\mathbb{F}_{p}^{m}$ by $\Phi(\mathbf{s}):=(f(\phi(\mathbf{s})_{1}),\ldots,f(\phi(\mathbf{s})_{m}))$. We refer to $\Phi(\mathbf{s})$ as the check vector of $\mathbf{s}$.
Let $H$ be a $2t\times m$ parity-check matrix of an (m,m2t)(m,m-2t) GRS code over $\mathbb{F}_{p}$. We now generate the syndrome of the check vector of $\bar{\mathbf{x}}$ with respect to $H$ and represent it as a $q$-ary string over $\Sigma_{q}$. Note that the syndrome $H\Phi(\bar{\mathbf{x}})^{\top}$ is a length-$(2t)$ vector over $\mathbb{F}_{p}$. Fix an injective mapping $g\colon\mathbb{F}_{p}\to\Sigma_{q}^{\lceil\log_{q}p\rceil}$. Let $\mathbf{v}\in\Sigma_{q}^{2t\lceil\log_{q}p\rceil}$ be the $q$-ary string obtained by applying $g$ to each entry of $H\Phi(\bar{\mathbf{x}})^{\top}$ and concatenating the results.
We next protect the $q$-ary string $\mathbf{v}$ against $t$ deletions. To do so, we encode $\mathbf{v}$ using a $q$-ary $t$-deletion-correcting code, for example, one from [36]. Denote the resulting encoding of $\mathbf{v}$ by $\mathscr{R}(\bar{\mathbf{x}})=\mathscr{R}(\mathbf{s}_{0},\mathbf{u})$. Finally, given the adapter sequences $\mathbf{s}_{0}$ and $\mathbf{s}_{n+\ell}$, our $t$-deletion-correcting code $\mathscr{C}$ of length $n$ over $\Sigma_{q}$ for the nanopore channel is defined to be
$\mathscr{C}=\{(\mathbf{u},\mathscr{R}(\mathbf{s}_{0},\mathbf{u}))\mid\mathbf{u}\in\Sigma_{q}^{k}\}.$
Remark III.1.
Note that the length of the $q$-ary string $\mathbf{v}$ representing the syndrome of the check vector for $\bar{\mathbf{x}}$ is $2t\lceil\log_{q}p\rceil\leq 2t\lceil\log_{q}(2k+4)\rceil$, since there always exists at least one prime number between $k+2$ and $2k+4$. If $q\leq 2t\lceil\log_{q}p\rceil$, which is typically the case for DNA sequencing, then by the results of [36], the length of $\mathscr{R}(\bar{\mathbf{x}})$ is $2t\log_{q}k+\Theta(\log\log k)$, assuming $k$ grows while $t,q$ are fixed constants independent of $k$. Therefore, the redundancy of the code $\mathscr{C}$ is $2t\log_{q}n+\Theta(\log\log n)$.
III-B Decoding
In this subsection, we show that any codeword $\mathbf{x}\in\mathscr{C}$ can be recovered from the output $\ell$-mers of the nanopore channel, provided that $\tilde{\mathbf{x}}=(\mathbf{s}_{0},\mathbf{x},\mathbf{s}_{n+\ell})$ is the input and at most $t\leq\{\ell-2,(\ell+1)/2\}$ deletions occur in the output. As before, let $\tau\leq t$ be the number of deletions imposed by the nanopore channel, which can be determined by counting the number of output $\ell$-mers. We assume the output $\ell$-mers $\mathbf{z}_{0}^{{n+\ell}-\tau}$ are consistent and $\tilde{\mathbf{y}}=\psi(\mathbf{z}_{0}^{{n+\ell}-\tau})$ is the corresponding output string. Moreover, let us write $\tilde{\mathbf{y}}=(y_{-\ell+1},\dots,y_{{n+\ell}-\tau})$. Denote the length of $\mathscr{R}(\bar{\mathbf{x}})=\mathscr{R}(\mathbf{s}_{0},\mathbf{u})$ by $r$, so the length of $\tilde{\mathbf{x}}$ is $n+2\ell=k+r+2\ell$.
The first step of our decoding procedure is to recover the syndrome $H\Phi(\bar{\mathbf{x}})^{\top}$ over $\mathbb{F}_{p}$. Since $\tau$ symbols of $\tilde{\mathbf{x}}$ are deleted, $\mathscr{R}(\bar{\mathbf{x}})$ can get shifted in $\tilde{\mathbf{y}}$ or deleted by at most $\tau$ positions. Therefore, $(y_{k+1},\dots,y_{k+r-\tau})$ must be a subsequence of $\mathscr{R}(\bar{\mathbf{x}})$. Since $\mathscr{R}(\bar{\mathbf{x}})$ is a codeword in a $q$-ary $t$-deletion-correcting code, we can recover $\mathscr{R}(\bar{\mathbf{x}})$ from $(y_{k+1},\dots,y_{k+r-\tau})$ and further obtain the $q$-ary string $\mathbf{v}$ that represents the syndrome of the check vector $\Phi(\bar{\mathbf{x}})$ with respect to $H$. Applying the mapping $g^{-1}$ to every $\lceil\log_{q}p\rceil$ symbols of $\mathbf{v}$, we recover the syndrome $H\Phi(\bar{\mathbf{x}})^{\top}$.
[…]
It remains to show that $\bar{\mathbf{x}}$ can be recovered given $\tilde{\mathbf{y}}$ and $\phi(\bar{\mathbf{x}})$. If $\hat{\mathbf{y}}$ is exactly $\bar{\mathbf{y}}$, then by Lemma 6 and 4, $\bar{\mathbf{x}}$ can be found by simply extending each check pattern in $\hat{\mathbf{y}}$ so that its length matches that of its counterpart in $\bar{\mathbf{x}}$. However, $\hat{\mathbf{y}}$ is generally only a prefix of $\bar{\mathbf{y}}$. So a careful analysis is needed to distinguish different cases.
[…]
Consider the case where $m_{\hat{\mathbf{y}}}=m_{\bar{\mathbf{x}}}$. In this case, the last check pattern in $\bar{\mathbf{x}}$ may remain intact or get shortened in the $\hat{\mathbf{y}}$. Therefore, we can obtain a string $\hat{\mathbf{x}}$ from $\hat{\mathbf{y}}$ and $\phi(\bar{\mathbf{x}})$ by extending the $i$-th check pattern in $\hat{\mathbf{y}}$ to be of length $\phi(\bar{\mathbf{x}})_{i}$ using its periodicity for all $i=1,\ldots,m_{\hat{\mathbf{y}}}$. Since $\hat{\mathbf{y}}$ is a prefix of $\bar{\mathbf{y}}$ obtained by truncating $\bar{\mathbf{y}}$ to its first $\ell+k-\tau$ symbols, the length of $\hat{\mathbf{x}}$ is no larger than that of $\bar{\mathbf{x}}$. Denote the length $\hat{\mathbf{x}}$ by $\ell+\hat{k}$. If $\hat{k}=k$, then the truncation occurs right at the ending position of $\bar{\mathbf{y}}$, and thus we have $\hat{\mathbf{x}}=\bar{\mathbf{x}}$.
If $\hat{k}<k$, then some symbols in $\bar{\mathbf{y}}$ that follow its last check pattern are removed when forming $\hat{\mathbf{y}}$. In fact, some symbols of the last check pattern in $\bar{\mathbf{y}}$ may also be removed by the truncation. However, in that case these symbols are already recovered in $\hat{\mathbf{x}}$ by extending the last check pattern in $\hat{\mathbf{y}}$ to have length $\phi(\bar{\mathbf{x}})_{m_{\bar{\mathbf{x}}}}$. At the same time, these symbols still appear in the substring $(y_{k+1-\tau},\ldots,y_{n+\ell-\tau})$ that follows $\hat{\mathbf{y}}$. Thus, we need to distinguish these two cases.
Let $\tau^{*}\leq\tau$ be the number of symbol deletions in $\bar{\mathbf{x}}$ induced by the nanopore channel to form $\bar{\mathbf{y}}$. Then the length difference between $\bar{\mathbf{y}}$ and $\hat{\mathbf{y}}$ is $\tau-\tau^{*}$ and we have $k-\hat{k}\leq\tau-\tau^{*}$. If the last check pattern in $\hat{\mathbf{y}}$ stops before the last symbol in $\hat{\mathbf{y}}$, then no symbols of the last check pattern in $\bar{\mathbf{y}}$ are removed by the truncation. It follows that $k-\hat{k}=\tau-\tau^{*}$ and we have $(\hat{\mathbf{x}},y_{k+1-\tau},\ldots,y_{2k-\hat{k}-\tau})=\bar{\mathbf{x}}$. If the last check pattern in $\hat{\mathbf{y}}$ stops at the last symbol in $\hat{\mathbf{y}}$, then we need to find the prefix of $(y_{k+1-\tau},\ldots,y_{n+\ell-\tau})$ that, when appended to $\hat{\mathbf{y}}$, extends its last check pattern in $\hat{\mathbf{y}}$. Note that such a prefix can be easily found by exploiting the periodicity of the last check pattern in $\hat{\mathbf{y}}$. Let $k+s-\tau$ be the ending index of this prefix where $s\geq 1$. Then $s+k-\hat{k}=\tau-\tau^{*}$ and thus we have $(\hat{\mathbf{x}},y_{k+s+1-\tau},\ldots,y_{2k+s-\hat{k}-\tau})=\bar{\mathbf{x}}$.
Lastly, consider the case where $m_{\hat{\mathbf{y}}}=m_{\bar{\mathbf{x}}}-1$. In this case, the last check pattern in $\bar{\mathbf{y}}$ is destroyed by truncating $\bar{\mathbf{y}}$ to $\hat{\mathbf{y}}$. More precisely, the truncation reduces its length to less than $\ell-1$. By definition of $m_{\hat{\mathbf{y}}}$, at most the last $\ell-2$ symbols of $\hat{\mathbf{y}}$ (and hence $\hat{\mathbf{x}}$) can belong to this check pattern. To identify it, we search for the first check pattern in $(y_{k-\tau-\ell+3},\ldots,y_{n+\ell-\tau})$. We then extend the last $\ell-2$ symbols of $\hat{\mathbf{x}}$ up to the first check pattern in $(y_{k-\tau-\ell+3},\ldots,y_{n+\ell-\tau})$, and subsequently extend the check pattern to have length $\phi(\bar{\mathbf{x}})_{m_{\bar{\mathbf{x}}}}$ using its periodicity. Note that the resulting string still has length at most $\ell+k$. By the same argument as in the case $m_{\hat{\mathbf{y}}}=m_{\bar{\mathbf{x}}}$, the resulting string either has length exactly $\ell+k$, in which case it coincides with $\bar{\mathbf{x}}$, or it can be further extended to obtain $\bar{\mathbf{x}}$.
After recovering $\bar{\mathbf{x}}$, the string $\mathbf{x}$ can be found by removing the adapter sequence $\mathbf{s}_{0}$. This leads to the following theorem.
Theorem 8.
Assume $t\leq\min\{\ell-2,(\ell+1)/2\}$. The code $\mathscr{C}\subset\Sigma_{q}^{n}$ described in Section III-A is a $t$-deletion-correcting code for the nanopore channel with redundancy $2t\log_{q}n+\Theta(\log\log n)$.
IV Existential bounds on the optimal code size
In this section, we present bounds on the maximum size of $t$-deletion-correcting codes for the nanopore channel. We begin with a simple observation: by Proposition 2, any $t$-deletion-correcting code for the $q$-ary deletion channel (i.e., a channel that takes strings over $\Sigma_{q}$ as input and deletes at most $t$ symbols) can also be used for the nanopore channel. This is because the adapter sequences are known and, as mentioned in Remark II.3, the nanopore channel, in effect, only deletes symbols within consistency-preserving patterns of its input, while the $q$-ary deletion channel is not subject to this restriction. Thus, any lower bound on the maximum size of codes for the $q$-ary deletion channel is also a lower bound for the nanopore channel. To the best of our knowledge, the best such bound (for the parameter regime considered in this work) is due to Alon et al. [37], yielding the following lower bound on the maximum size of codes for the nanopore channel.
Theorem 9 ([37, Theorem 1]).
A $t$-deletion-correcting code $\mathscr{C}\subset\Sigma_{q}^{n}$ for the nanopore channel has size $|\mathscr{C}|=\Omega(q^{n}n^{-2t}\log n)$.
Remark IV.1.
By Theorem 9, there exist $t$-deletion-correcting codes of length $n$ over $\Sigma_{q}$ for the nanopore channel with redundancy at most $2t\log_{q}n-\log_{q}\log_{2}n+O(1)$. However, this result is nonconstructive. In contrast, the explicit codes presented in Section III have redundancy $2t\log_{q}n+\Theta(\log\log n)$, which matches the bound in Theorem 9 to first order.
Next, we derive an upper bound on the size of $t$-deletion-correcting codes for the nanopore channel. For convenience, a string with period one is said to be constant. Our approach is to consider a relaxed setting in which the nanopore channel is restricted to deleting symbols only within sufficiently long maximal constant substrings of the input. The largest possible size of a $t$-deletion-correcting code in this relaxed setting then serves as an upper bound on the maximum code size in the original setting. Intuitively, a code for this relaxed setting cannot contain many codewords that have a large number of maximal constant substrings. Motivated by this observation, we adapt the ideas of Levenshtein [29] and use the probabilistic method. Specifically, we view each symbol of the string to be sequenced as i.i.d. and uniformly distributed over $\Sigma_{q}$, and estimate the number of strings with a small number of maximal constant substrings from this perspective. Furthermore, since the symbols within a maximal constant substring are not independent, we employ Janson’s concentration inequality [45] for dependent random variables to bound the number of such strings. Below we state Janson’s concentration inequality and then present our upper bound.
Theorem 10 (Janson’s concentration inequality [45, Theorem 2.1]).
Let $X=\sum_{\alpha\in A}Y_{\alpha}$ where $Y_{\alpha}$ is a random variable with $a_{\alpha}\leq Y_{\alpha}\leq b_{\alpha}$ and some real numbers $a_{\alpha}$ and $b_{\alpha}$. Then for $\eta>0$,
$\Pr(X\geq\mathbb{E}X+\eta)\leq\exp\Big(-2\frac{\eta^{2}}{\chi^{*}(\Gamma)\sum_{\alpha\in A}(b_{\alpha}-a_{\alpha})^{2}}\Big),$
where $\Gamma$ is a dependency graph for $\{Y_{\alpha}\}_{\alpha\in A}$ and $\chi^{*}(\Gamma)$ is the fractional chromatic number of $\Gamma$. The same estimate holds for $\Pr(X\leq\mathbb{E}X-\eta)$.
Theorem 11.
A $t$-deletion-correcting code $\mathscr{C}\subset\Sigma_{q}^{n}$ for the nanopore channel has size $|\mathscr{C}|=O(q^{n}n^{-t})$.
Proof.
Consider a weaker version of the adversarial nanopore channel where deletions only occur within consistency-preserving patterns of $\mathbf{x}$ that are constant substrings, i.e., within maximal constant substrings of length at least $\ell$ in $\mathbf{x}$. Denote by $M$ the largest possible size of a $t$-deletion-correcting codes for this weaker channel. Clearly, $|\mathscr{C}|\leq M$. In the following, we derive an upper bound on $M$. Our derivation is inspired by [29, Lemma 2].
Let $\mathscr{C}^{*}\subset\Sigma_{q}^{n}$ be a $t$-deletion-correcting code for the weaker channel with $|\mathscr{C}^{*}|=M$. Partition $\mathscr{C}^{*}$ into two subsets $\mathscr{C}_{1},\mathscr{C}_{2}$ according to the number of maximal constant substrings in each of the codewords. More precisely, for an string $\mathbf{x}\in\Sigma_{q}^{n}$, let $\lVert\mathbf{x}\rVert$ be the number of maximal constant substrings in $\mathbf{x}$. For some integer $\kappa\geq t$, define
$\mathscr{C}_{1}=\{\mathbf{x}\in\mathscr{C}^{*}:\lVert\mathbf{x}\rVert>\kappa\},\quad\mathscr{C}_{2}=\{\mathbf{x}\in\mathscr{C}^{*}:\lVert\mathbf{x}\rVert\leq\kappa\}.$
Let $\mathscr{S}_{t}(\mathbf{x})$ be the set of all strings of length ntn-t that can be obtained from $\mathbf{x}$ by deleting symbols in maximal constant substrings in $\mathbf{x}$. Since $\mathscr{S}_{t}(\mathbf{x}),\mathbf{x}\in\mathscr{C}_{1}$ are disjoint, we have $\sum_{\mathbf{x}\in\mathscr{C}_{1}}\left|\mathscr{S}_{t}(\mathbf{x})\right|\leq q^{n-t}$. Moreover, since $\kappa\geq t$, $\left|\mathscr{S}_{t}(\mathbf{x})\right|\geq\binom{\kappa}{t}>0$. Therefore, we have $|\mathscr{C}_{1}|\leq q^{n-t}/\binom{\kappa}{t}$. In the following, we set $\kappa=\mu-\sqrt{(\ell tn\ln n)/2}$ for sufficiently large $n$ (so that $\kappa\geq t$), where
$\mu:=\frac{1+(n-\ell+1)(q-1)}{q^{\ell}}.$
As a consequence, we have $|\mathscr{C}_{1}|\leq q^{n-t}/\binom{\kappa}{t}=\Theta(q^{n}n^{-t})$.
We next bound the size of $\mathscr{C}_{2}$. If $\mathscr{C}_{2}$ also has size at most $\Theta(q^{n}n^{-t})$, then $|\mathscr{C}^{*}|\leq\Theta(q^{n}n^{-t})$. Let $\mathscr{D}:=\{\mathbf{x}\in\Sigma_{q}^{n}:\lVert\mathbf{x}\rVert\leq\kappa\}$. Since $\mathscr{C}_{2}\subset\mathscr{D}$, we have $|\mathscr{C}_{2}|\leq|\mathscr{D}|$. We next use a probabilistic approach to show that $|\mathscr{D}|\leq\Theta(q^{n}n^{-t})$. Write $\mathbf{x}=(x_{1},\ldots,x_{n})$. Assume that $x_{1},\ldots,x_{n}$ are i.i.d. random variables, each uniformly distributed on $\Sigma_{q}$. Then $|\mathscr{D}|=q^{n}\Pr(\mathbf{x}\in\mathscr{D})$. For $\mathbf{x}\in\Sigma_{q}^{n}$, define a sequence of indicator random variables $(I_{i})_{1\leq i\leq n-\ell+1}$ such that $I_{i}=1$ if and only if a maximal constant substring of length at least $\ell$ starts at the $i$-th position of $\mathbf{x}$. Then we have $\Pr(\mathbf{x}\in\mathscr{D})=\Pr(\lVert\mathbf{x}\rVert\leq\kappa)=\Pr\big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\kappa\big)$.
In the following, we bound $\Pr\big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\kappa\big)$ from above using Janson’s concentration inequality [45]. To this end, let us compute the expectation of $\sum_{i=1}^{n-\ell+1}I_{i}$. Note that $\Pr(I_{1}=1)=\Pr(x_{1+j}=x_{1}\text{ for }j=1,\dots,\ell-1)=\prod_{i=2}^{\ell}\Pr(x_{i}=x_{1})={q^{-\ell+1}}$. Moreover, for $2\leq i\leq n-\ell+1$ we have $\Pr(I_{i}=1)=\Pr(x_{i+j}=x_{i}\text{ for }j=1,\dots,\ell-1\text{ and }x_{i}\neq x_{i-1})=\Pr(x_{i}\neq x_{i-1})\prod_{i=2}^{\ell}\Pr(x_{i}=x_{1})=(q-1)/{q^{\ell}}$. Therefore, $\mathbb{E}\big[\sum_{i=1}^{n-\ell+1}I_{i}\big]=\sum_{i=1}^{n-\ell+1}\mathbb{E}[I_{i}]=\mu$.
Note that $I_{i}$ and $I_{j}$ are independent if and only if $|i-j|\geq\ell$. Let $\mathscr{G}=(V,E)$ be the interval graph on $n-\ell+1$ vertices where vertex $v_{i}\in V$ corresponds to closed interval $i,i+\ell-1]$ and $(v_{i},v_{j})\in E$ if and only if the corresponding intervals have a non-empty intersection. Therefore, $\mathscr{G}$ is perfect and its chromatic number equals its clique number, i.e., $\chi(\mathscr{G})=\omega(\mathscr{G})$ (e.g., see [46]). Since any $\ell$ consecutive vertices $v_{i},\dots,v_{i+\ell-1}$ form a clique while any set of more than $\ell$ vertices must include two vertices that are independent and not connected by an edge, we have $\omega(\mathscr{G})=\ell$. Note that the fractional chromatic number $\chi^{*}$ of a graph is bounded between its clique number $\omega$ and chromatic number $\chi$. It follows that $\omega(\mathscr{G})\leq\chi^{*}(\mathscr{G})\leq\chi(\mathscr{G})=\omega(\mathscr{G})$, and therefore, $\chi^{*}(\mathscr{G})=\ell$. Since $\mathscr{G}$ is a dependency graph for $(I_{i})_{1\leq i\leq n-\ell+1}$, by Theorem 10 we have
$\Pr\Big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\kappa\Big)$ $=\Pr\Big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\mu-\sqrt{(\ell tn\ln n)/2}\Big)$
$=\Pr\Big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\mathbb{E}\Big[\sum_{i=1}^{n-\ell+1}I_{i}\Big]-\sqrt{(\ell tn\ln n)/2}\Big)$
$\leq\exp\Big(-\frac{tn\ln n}{n-\ell+1}\Big)$
$\leq\exp(-t\ln n).$
Hence, $|\mathscr{D}|=q^{n}\Pr\big(\sum_{i=1}^{n-\ell+1}I_{i}\leq\kappa\big)\leq\Theta(q^{n}n^{-t})$. It follows that $|\mathscr{C}_{2}|\leq\Theta(q^{n}n^{-t})$. ∎
Remark IV.2.
By Theorem 11, the redundancy of $t$-deletion-correcting codes for the nanopore channel is at least $t\log_{q}n+\Omega(1)$.
V Concluding remarks
We note that the reduction from deletions to Hamming-type errors relies on the assumption that $t\leq\min\{\ell-2,(\ell+1)/2\}$. It is therefore natural to ask what can be done if $t$ is larger. Moreover, since nanopore sequencing errors also include duplications and substitutions, it would be interesting to further study codes for an adversarial nanopore channel that can introduce deletions, duplications and substitutions.